# The optimal hypercontractive constants for $\mathbb{Z}_{3}$

### v1

@@ 1.1. Main results
We believe that, the methods developed in this paper can also be used to establish the optimal hypercontractivity constant for the asymmetric two-point inequalities:
(1.7) $\|1+rx\varepsilon_{\alpha}\|_{q}\leq\|1+x\varepsilon_{\alpha}\|_{p}\quad\text{with}\quad\mathbb{P}(\varepsilon_{\alpha}=1-\alpha)=1-\mathbb{P}(\varepsilon_{\alpha}=-\alpha)=\alpha.$
Wolff [Wol07, Theorems 2.1 and 2.2] proved that the optimal hypercontractivity constant $\sigma_{p,q}(\alpha)$ for (1.7) is equivalent, up to a universal numerical factor, to some explicit constant $\tilde{\sigma}_{p,q}(\alpha)$. Moreover, he [Wol07, Theorem 3.1] proved that the hypercontractivity constant for any finite discrete measure is controlled by (1.7) with $\alpha$ being the mass of its smallest atom. In particular, for $p=2$, this result combined with Latała and Oleszkiewicz [LaO00, Ole03] allows Wolff to compute $r_{2,q}(\mathbb{Z}_{3})$.
@@ 1.2.1. Reducing the full space to a segment
The reduction proceeds in two main steps, each narrowing the candidate space for non-trivial extremizers. While the first reduction is relatively routine — using standard symmetry arguments — the second reduction is technical and crucial. It effectively lowers the dimension of the parameter space from $\mathbb{C}^{3}$ to a one-dimensional interval $[0,1]$. This dimensional reduction is what ultimately allows us to formulate and solve a tractable system of equations for the extremal parameters in the subsequent analysis.
@@
Reduction 2: from $\Delta$ to $[0,1]$ via monotonicity. For any $1<p<q<\infty$, noting that by duality, $r_{p,q}=r_{q^{*},p^{*}}$, it suffices to determine $r_{p,q}$ for the cases
@@ 1.2.2. Establishing the system of equations
Indeed, since $\rho_{0}\in(0,1)$ is an interior point of $[0,1]$ where the function $G(r_{p,q},\cdot)$ attains its minimum $G(r_{p,q},\rho_{0})=0$, the pair $(r_{p,q},\rho_{0})$ satisfies the system
@@
The remainder of the proof is devoted to establishing that (1.2) admits a unique solution $(x,y)$ in the interior domain $(0,1)\times(0,1)$.
@@ 1.2.3. Uniqueness of the solution
Blowup analysis. To rigorously prove the unique intersection property of the new curve family $H(\alpha,\cdot)$, we require three properties: injectivity, separation, and transversality, where the latter two are to be established via appropriate blowups.
@@
Organization of the paper: The remaining part of the paper is organized as follows. In Section 2, we present the orthogonal basis method and the monotonicity game, which are fundamental to our subsequent analysis. Section 3 is devoted to reducing the domain of the function coefficients in the hypercontractive inequalities from $\mathbb{C}^{3}$ to the interval $[0,1]$. In Section 4, we first prove the existence of nontrivial extremizers for the $\mathbb{Z}_{3}$ hypercontractive inequality and then, by applying the variational principle, obtain a characterizing system of equations. The proof of the main theorem is given in Section 5, contingent on the uniqueness of the solution to this system, which is addressed in Section 6. Lastly, the appendix provides elementary proofs for several inequalities and for the corollaries of the main result.
Acknowledgements. JC is supported by Postdoctoral Daily Funding (No.000141020162),SF is supported by NSFC (No.12331004 and No.12231013), YH is supported by NSFC (No.12131016), ZW is supported by NSFC (No.12471116) and FRFCU (No.2025CDJ-IAIS YB004).
@@ 2.1. Orthogonal basis method
Let $A$ be a strictly convex or strictly concave function on an interval $(a,b)\subset\mathbb{R}$, and let $(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}$ be a nonzero vector such that $y_{1}+y_{2}+y_{3}=0$. If $x_{1},x_{2},x_{3}\in(a,b)$ are distinct and satisfy $x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}=0$, then
@@
We proceed by contradiction. Assume that
$\sum_{k=1}^{3}A(x_{k})y_{k}=0.$
That is, the vector $(A(x_{1}),A(x_{2}),A(x_{3}))$ is orthogonal to the nonzero vector $(y_{1},y_{2},y_{3})$. By assumption, the vectors $(1,1,1)$ and $(x_{1},x_{2},x_{3})$ are also orthogonal to $(y_{1},y_{2},y_{3})$. Moreover, the vectors $(1,1,1)$ and $(x_{1},x_{2},x_{3})$ are linearly independent since $x_{1},x_{2},x_{3}$ are distinct. Thus there exist real numbers $s,t\in\mathbb{R}$ such that
@@
In our proofs of the monotonicity lemmas, the relevant setting satisfies a stronger condition: the three vectors
@@ 3. Parameter reduction: from $\mathbb{C}^{3}$ to $[0,1]$
Let $1<p<q<\infty$. The hypercontractivity problem for $\mathbb{Z}_{3}$ is to determine the optimal constant $r_{p,q}=r_{p,q}(\mathbb{Z}_{3})$ such that for $r\in(0,r_{p,q}]$,
@@
Note that $0<r_{p,q}<\sqrt{(p-1)/(q-1)}$ seems well-known. For completeness, we state it as follows.
@@
Let $1<p<q<2$ or $1<p\leq 2<q<\infty$. Then for any $r\in[0,\sqrt{(p-1)/(q-1)}]$, the inequality (3.1) holds for all functions on $\mathbb{Z}_{3}$ if and only if it holds for all functions of the form
@@
Clearly, the coefficient of $a^{3}$ is nonzero. Therefore, we get $\|f_{a}\|_{p}-\|T_{r}f_{a}\|_{q}<0$ for some $a$ near 0. It follows that $r_{p,q}<\sqrt{(p-1)/(q-1)}$, as desired.
@@ 3.2. Reduction from $\mathbb{C}^{3}$ to $\Delta$ via symmetry
The Fourier expansion of any function $f$ on $\mathbb{Z}_{3}$ is given by (noting that $\chi^{2}$ coincides with $\overline{\chi}$)
@@ 3.3. Reduction from $\Delta$ to $[0,1]$ via monotonicity
We begin by noting that triangle $\Delta$ can be expressed in polar coordinates by
@@
We are now ready to prove the monotonicity of the defect function $G$ in $\theta$ for the case $1<p\leq 2<q<\infty$ by combining the derivative estimates above.
@@
Fix $\rho\in(0,1)$. Note that $F(2,\rho,\theta)=3(1+2\rho^{2})$ is constant with respect to $\theta$. For $1<p\leq 2<q<\infty$, Lemma 3.7 implies that $F(p,\rho,\theta)$ is increasing in $\theta$, while $F(q,r\rho,\theta)$ is strictly decreasing in $\theta$. It follows that the defect function
@@
In [Bor82], Borell established a reverse hypercontractive inequality on $\mathbb{Z}_{2}$. Recently, Mossel, Oleszkiewicz, and Sen [MOS13] showed that reverse hypercontractive inequalities follow from standard hypercontractive inequalities or the modified log-Sobolev inequality. In particular, they obtained that Borell’s reverse hypercontractive inequalities hold uniformly for all probability spaces.
@@ 4. Establishing the equation-system via variational principle
By Proposition 3.2, the $(p,q)$-hypercontractivity problem for $\mathbb{Z}_{3}$ is reduced to finding the optimal $r_{p,q}(\mathbb{Z}_{3})\in[0,1]$ such that
$\left\|T_{r}f_{\rho}\right\|_{q}\leq\|f_{\rho}\|_{p}\quad\text{for all $f_{\rho}=1+\rho\chi+\rho\overline{\chi}$ with $\rho\in[0,1]$.}$
@@
  • establishing the existence of a nontrivial extremizer,
@@
4.1. Existence of a nontrivial extremizer

### v2

@@ 1.2. Optimal hypercontractive constants for biased Bernoulli random variables
The reader may notice that, in Theorem 1.6, we do not state the uniqueness of the solution in the unit square of the equation-system (1.10). Indeed, with a few modifications, the proof of the uniqueness statement in the setting of Theorem 1.6 follows almost verbatim from that of Theorem 1.2 for large $\lambda$ (say $1/5\leq\lambda<1/2$). However, for smaller parameter $\lambda$, some additional work is needed. Therefore, in this general setting of Theorem 1.6, we have decided to include all the proofs of the uniqueness in a separate document not intended for publication.
Remark.
For small parameter $\lambda$, a brief description of the scenario is given below. For each $\lambda\in(0,1/2)$, define
$\begin{array}[]{rccc}h_{\lambda}:&(1,\infty)\times(0,1)&\longrightarrow&\mathbb{R}\times\mathbb{R}\\ &(p,x)&\longmapsto&\Big(\frac{(\lambda+(1-\lambda)x^{p})^{1/p}}{\lambda+(1-\lambda)x},\;\frac{(1-x)(1-x^{p-1})}{\lambda+(1-\lambda)x^{p}}\Big).\end{array}$
Since 1x1-x, $1-x^{p-1}$, $1/(\lambda+(1-\lambda)x^{p})$ are all strictly decreasing in $x$, the ordinate of the curve $h_{\lambda}(p,\cdot)$ decreases in $x$. So $h_{\lambda}(p,\cdot)$ has no self-intersecting point. It follows that the system (1.10) admits a unique solution in $(0,1)^{2}$ if and only if the curves $h_{\lambda}(p,\cdot)$ and $h_{\lambda}(q,\cdot)$ intersect exactly once. Through suitable changes of coordinates, we will obtain another curve family
$\begin{array}[]{rccc}H_{\lambda}=(H_{\lambda,1},H_{\lambda,2}):&(-1,1)\times(0,1)&\longrightarrow&\mathbb{R}\times\mathbb{R}\\ &(\alpha,t)&\longmapsto&\left(\begin{array}[]{c}\tfrac{\alpha\log(\lambda+(1-\lambda)t^{2})-\log(\lambda+(1-\lambda)t^{1+\alpha})+\log(\lambda+(1-\lambda)t^{1-\alpha})}{\log\lambda}\\[5.69054pt] \tfrac{-\log(\lambda+(1-\lambda)t^{2})+\log(\lambda+(1-\lambda)t^{1+\alpha})+\log(\lambda+(1-\lambda)t^{1-\alpha})}{\log\lambda}\end{array}\right).\end{array}$
Direct computation gives the following properties.
  • Symmetries $H_{\lambda,1}(-\alpha,t)=-H_{\lambda,1}(\alpha,t)$ and $H_{\lambda,2}(-\alpha,t)=H_{\lambda,2}(\alpha,t)$.
  • The end points of $H_{\lambda}(\alpha,\cdot)$ are $H_{\lambda}(\alpha,0)=(\alpha,1)$ and $H_{\lambda}(\alpha,1)=(0,0)$.
  • Since $H_{\lambda,1}(0,t)\equiv 0$, we have $H_{\lambda}(0,[0,1])=\{0\}\times[0,1]$.
  • Since $H_{\lambda,1}(\alpha,\tfrac{\lambda}{1-\lambda})\equiv 0$, we have $H_{\lambda}(\alpha,\tfrac{\lambda}{1-\lambda})\in\{0\}\times[0,1]$.
Furthermore, we have the following observations from graphs.
  • For $\alpha\neq\beta$, the curves $H_{\lambda}(\alpha,\cdot)$ and $H_{\lambda}(\beta,\cdot)$ intersect exactly once in the upper half-plane. This is the evidence that the system (1.10) admits a unique solution in $(0,1)^{2}$.
  • The singularity $(\alpha,t)=(0,\tfrac{\lambda}{1-\lambda})$ is a Whitney pleat and
    $H_{\lambda}(0,\tfrac{\lambda}{1-\lambda})=(0,\tfrac{\log(4\lambda(1-\lambda))}{\log\lambda})\xrightarrow[\lambda\to 0]{}(0,1).$
  • When $\lambda\to 0$, the curve family $H_{\lambda}$ has a limit shape as shown in Figure 2.
Figure 2. The curve family $\{H_{\lambda}(\alpha,\cdot)\}_{\alpha\in(-1,1)}$ with $\lambda=10^{-100}$
Remark.
Similar to (1.5), one also has a description of $\sigma_{p,q}(\lambda)$ that avoids relying on the uniqueness of the solution in $(0,1)^{2}$ to (1.10):
[…]
Let $\mu$ be a probability measure on a finite set $\Omega$, for any $1<p<q<\infty$, define
$\sigma_{p,q}(\mu):$ $=\max\big\{0<r<1:\|\mathbb{E}_{\mu}+r(Id-\mathbb{E}_{\mu})\|_{L^{p}(\mu)\rightarrow L^{q}(\mu)}\leq 1\big\}$
[…]
Since $\sigma_{p,p^{*}}(\lambda)$ given in (1.11) is increasing in $\lambda\in(0,1/2)$, using [Wol07, Theorem 3.1], we have
Corollary 1.7.
Let $\mu$ be a probability measure on a finite set $\Omega$ with the mass of the least atom equal to $\lambda\in(0,1/2)$. Then for all $1<p<2$, we have $\sigma_{p,p^{*}}(\mu)=\sigma_{p,p^{*}}(\lambda)$ which is given by (1.11).
Remark.
Since $\sigma_{p,q}(\lambda)=\sigma_{p,q}(1-\lambda)$, Theorem 1.6 gives also $\sigma_{p,q}(\lambda)$ for all $\lambda\in(0,1)\setminus\{1/2\}$.
Remark.
Numerical simulation indicates that $\sigma_{p,q}(\lambda)$ is increasing in $\lambda\in(0,1/2]$.
Figure 3. Numerical simulation of $\sigma_{p,q}(\lambda)$
Inspired by the simulations in Figure 3, using [Wol07, Theorem 3.1], we formulate the following
Conjecture.
Let $\mu$ be a probability measure on a finite set $\Omega$ with the mass of the least atom equal to $\lambda\in(0,1/2)$. Then for all $1<p<q<\infty$, we have $\sigma_{p,q}(\mu)=\sigma_{p,q}(\lambda)$.
1.3. Noise operators
The noise operator plays an important role in $\lambda$-biased Fourier analysis, and provides another key example of hypercontractive random variables. For $\lambda\in(0,1/2]$, let $\Omega=\{-1,1\}$ and let $\pi_{\lambda}$ be the distribution on $\Omega$ defined by
$\pi_{\lambda}(-1)=\lambda\quad\text{and}\quad\pi_{\lambda}(1)=1-\lambda.$
An orthogonal basis (Fourier basis) for the real Hilbert space $L^{2}(\Omega,\pi_{\lambda})$ consists of the constant function $1$ and
$\phi_{\lambda}(x)=\frac{x-(1-2\lambda)}{2\sqrt{\lambda(1-\lambda)}},\quad x\in\Omega.$
For $r\in(0,1]$, the noise operator with parameter $r$ is the linear operator on $L^{2}(\Omega,\pi_{\lambda})$ defined by
(1.12) $T_{r}f=1+r\rho\phi_{\lambda},\quad\text{with}\quad f=1+\rho\phi_{\lambda}\quad\text{and}\quad\rho=\mathbb{E}_{x\sim\pi_{\lambda}}[f(x)\phi_{\lambda}(x)].$
In this setting, the hypercontractive random variable is given by $Y_{\lambda}=\phi_{\lambda}$:
(1.13) $\mathbb{P}\Big(Y_{\lambda}=\sqrt{\lambda/(1-\lambda)}\Big)=1-\lambda\text{\,\, and \,\,}\mathbb{P}\Big(Y_{\lambda}=-\sqrt{(1-\lambda)/\lambda}\Big)=\lambda.$
The related problem is known as $\lambda$-biased hypercontractive problem (see [O’D14, Chapters 8 and 9]). Further details can be found in [O’D14, Definitions 8.26, 8.27, 8.39 and Proposition 8.27].
By the works of Bourgain-Kahn-Kalai-Katznelson-Linial [BKKKL92], Talagrand [Tal94], Friedgut-Kalai [FK96], and Friedgut [Fri98], the general hypercontractive theorem [O’D14, Page 278] states that for every $f\in L^{2}(\Omega,\pi_{\lambda})$,
$\|T_{r}f\|_{q}\leq\|f\|_{2}\quad\text{whenever}\quad 0\leq r\leq\frac{1}{\sqrt{q-1}}\lambda^{\,1/2-1/q}.$
It is observed that when $\lambda$ is small, the hypercontractive inequality requires $r$ to also be small, and such a bound is often too weak for applications. In a recent breakthrough, Keevash, Lifshitz, Long, and Minzer [KLLM24] introduced global hypercontractivity, a new form of biased hypercontractive inequality. This tool yields a strengthened version of Bourgain’s theorem, advances on a conjecture of Kahn and Kalai, and a $\lambda$-biased analogue of the Mossel–O’Donnell–Oleszkiewicz invariance principle. Subsequently, Keller, Lifshitz, and Marcus [KLM26] established sharp global hypercontractivity, leading to numerous applications in extremal set theory, group theory, theoretical computer science, and number theory. From the concluding remarks of [KLLM24], global hypercontractivity yields a variant of the Kahn–Kalai isoperimetric conjecture that is effective only in the $\lambda$-biased setting for small $\lambda$; corresponding results for the uniform measure are given in [KLo18] and [KLi18]. The authors posed the open problem of finding a unified approach that extends both results to all $\lambda\in(0,1/2]$. However, we note that the correct bound for general hypercontractivity is unknown [Zha21].
One may notice that the hypercontractive analysis of the noise operator defined in (1.12) can be transferred to that of the random variable $Y_{\lambda}$ defined as (1.13). Indeed, one has
$Y_{\lambda}=-\frac{1}{\sqrt{\lambda(1-\lambda)}}X_{\lambda}$
and hence $Y_{\lambda}$ and $X_{\lambda}$ share the same optimal hypercontractive constant. Therefore, Theorem 1.6 provides the optimal hypercontractive constant for the noise operator defined in (1.12).
@@ 1.4.1. Reducing the full space to a segment
The reduction proceeds in two main steps, each narrowing the candidate space for nontrivial critical extremizers. While the first reduction is relatively routine — using standard symmetry arguments — the second reduction is technical and crucial. It effectively lowers the dimension of the parameter space from $\mathbb{C}^{3}$ to a one-dimensional interval $[0,1]$. This dimensional reduction is what ultimately allows us to formulate and solve a tractable system of equations for the extremal parameters in the subsequent analysis.
Remark.
An alternative way of reducing the space for candidate nontrivial critical extremizers is to use Wolff’s argument Wolff [Wol07, Theorem 3.1]: assuming their existence, they must be exactly two-valued. In this paper, we choose to give a self-contained direct reduction which seems to be of independent interest. Moreover, the proofs provide key strategies and insights for our proof of the uniqueness of the solution to (1.4). It can therefore be seen as a prelude to that proof.
@@
Reduction 2: from $\Delta$ to $[0,1]$ via monotonicity. By (1.6), it suffices to determine $r_{p,q}$ for the two cases
@@ 1.4.2. Establishing the system of equations
More precisely, since $\rho_{0}\in(0,1)$ is an interior point of $[0,1]$ where the function $G(r_{p,q},\cdot)$ attains its minimum $G(r_{p,q},\rho_{0})=0$, the pair $(r_{p,q},\rho_{0})$ satisfies the system
@@
The remainder of the proof is devoted to establishing that (1.4) admits a unique solution in the interior domain $(0,1)\times(0,1)$.
@@ 1.4.3. Uniqueness of the solution
Blowup analysis. To rigorously prove the unique intersection property of the new curve family $H(\alpha,\cdot)$, we require three properties: injectivity, separation, and transversality, where the latter two are established via appropriate blowups.
@@
Organization of the paper: The remaining part of the paper is organized as follows. In Section 2, we present the orthogonal basis method and the monotonicity game, which are fundamental to our subsequent analysis. Section 3 is devoted to reducing the domain of the function coefficients in the hypercontractive inequalities from $\mathbb{C}^{3}$ to the interval $[0,1]$. In Section 4, we first prove the existence of nontrivial critical extremizers for the $\mathbb{Z}_{3}$ hypercontractive inequality and then, by applying the variational principle, obtain a characterizing system of equations. The proof of the main theorem is given in Section 5, contingent on the uniqueness of the solution to this system, which is addressed in Section  6. Section 7 is devoted to a sketched proof of Theorem 1.6. Lastly, the appendix provides elementary proofs for several inequalities and for the corollaries of the main result.
Acknowledgements. We would like to thank Prof. Javier Parcet and Dr. Gan Yao for helpful discussions. JC is supported by Postdoctoral Daily Funding (No.000141020162), SF is supported by NSFC (No.12331004 and No.12231013), YH is supported by NSFC (No.12131016), YQ is supported by NSFC (No.12595283 and No.12471145), ZW is supported by NSFC (No.12471116) and FRFCU (No.2025CDJ-IAIS YB004).
@@ 2.1. Orthogonal basis method
Let $A$ be a strictly convex or strictly concave function on an interval $(a,b)\subset\mathbb{R}$, and let $(y_{1},y_{2},y_{3})\in\mathbb{R}^{3}$ be a nonzero vector with $y_{1}+y_{2}+y_{3}=0$. If $x_{1},x_{2},x_{3}\in(a,b)$ are distinct and satisfy $x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3}=0$, then
@@
Assume by contradiction that $\sum_{k=1}^{3}A(x_{k})y_{k}=0.$ That is, the vector $(A(x_{1}),A(x_{2}),A(x_{3}))$ is orthogonal to the nonzero vector $(y_{1},y_{2},y_{3})$. By assumption, the vectors $(1,1,1)$ and $(x_{1},x_{2},x_{3})$ are orthogonal to $(y_{1},y_{2},y_{3})$. And, the vectors $(1,1,1)$ and $(x_{1},x_{2},x_{3})$ are linearly independent since $x_{1},x_{2},x_{3}$ are distinct. Thus there exist $s,t\in\mathbb{R}$ such that
@@
In our proofs of the monotonicity lemmas, the relevant setting satisfies a stronger condition: the vectors
@@ 3. Parameter reduction: from $\mathbb{C}^{3}$ to $[0,1]$
Let $1<p<q<\infty$. Our goal is to determine the optimal $r_{p,q}$ such that for $r\in(0,r_{p,q}]$,
@@
It seems that $0<r_{p,q}<\sqrt{(p-1)/(q-1)}$ is well-known. For completeness, we state it as follows.
@@
Let $1<p<q<2$ or $1<p\leq 2<q<\infty$ and $r\in(0,\sqrt{(p-1)/(q-1)}]$. Then the inequality (3.1) holds for all functions on $\mathbb{Z}_{3}$ if and only if it holds for all functions of the form
@@
Since the coefficient of $a^{3}$ is nonzero, $\|f_{a}\|_{p}-\|T_{r}f_{a}\|_{q}<0$ for some small $a$. It follows that $r_{p,q}<\sqrt{(p-1)/(q-1)}$, as desired.
@@ 3.2. Reduction from $\mathbb{C}^{3}$ to $\Delta$ via symmetry
Any function $f$ on $\mathbb{Z}_{3}$ has Fourier expansion:
@@ 3.3. Reduction from $\Delta$ to $[0,1]$ via monotonicity
Write the triangle $\Delta$ in polar coordinates:
@@
We are now ready to prove the monotonicity of the defect function $G$ in $\theta$ for $1<p\leq 2<q<\infty$.
@@
Fix $\rho\in(0,1)$. For $1<p\leq 2<q<\infty$, Lemma 3.7 implies that $F(p,\rho,\theta)$ is increasing in $\theta$, while $F(q,r\rho,\theta)$ is strictly decreasing in $\theta$. It follows that the defect function
@@ 4. Establishing the equation-system via variational principle
Proposition 3.2 reduces the hypercontractivity for $\mathbb{Z}_{3}$ to finding the optimal $r_{p,q}(\mathbb{Z}_{3})$ such that
$\left\|T_{r}f_{\rho}\right\|_{q}\leq\|f_{\rho}\|_{p}\quad\text{for all $f_{\rho}=1+\rho\chi+\rho\overline{\chi}$ with $\rho\in[0,1]$ and all $0\leq r\leq r_{p,q}(\mathbb{Z}_{3})$.}$
@@
  • establishing the existence of a nontrivial critical extremizer,
@@
4.1. Existence of a nontrivial critical extremizer